OpenStudy (anonymous):

Write the equation of the line that is parallel to the line 3x − y = −3 and passes through the point (4, −2). Choose one answer. a. y = −one thirdx − 6 b. y = −one thirdx − 14 c. y = 3x − 6 d. y = 3x − 14

OpenStudy (anonymous):

so do the same as the first problem we had?

OpenStudy (anonymous):

You want the slope to stay the same the slop is 3x. So it has to be either c or d.

OpenStudy (anonymous):

then plug in x?

OpenStudy (sachintha):

Yeah same as before.

OpenStudy (anonymous):

its d

OpenStudy (anonymous):

i didnt need the answer. i needed help with solving it

OpenStudy (anonymous):

@sourav_aich don't just give away answers try to help people learn how to do the math so they can learn it.

OpenStudy (sachintha):

First turn the equation to the form y= mx+c

OpenStudy (anonymous):

the 3x- y = -3?

OpenStudy (sachintha):

yeah.

OpenStudy (anonymous):

how do i ove the 3x back?

OpenStudy (anonymous):

just sumbtract both sides?

OpenStudy (anonymous):

subtract*

OpenStudy (sachintha):

yeah and then divide the whole equation by -1 so you get positive y.

OpenStudy (anonymous):

y= 3x + 6

OpenStudy (anonymous):

+3 sorry not 6

OpenStudy (anonymous):

so y= 3x +3 What do i do from here

OpenStudy (sachintha):

Continue as you did before.

OpenStudy (anonymous):

so just plug the rest in y= 3x - 14 x=4 y= 12-14 y=-2

OpenStudy (anonymous):

so it i d

OpenStudy (anonymous):

I need help with one more so ill open another thread

OpenStudy (sachintha):

You are correct. :)